# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

### v3

@@ 6 Conclusion
In this study, we advanced Memory Gym’s environments to novel endless tasks, tailored to benchmark memory-based Deep Reinforcement Learning algorithms. These tasks feature a mounting challenge similar to the cumulative memory game “I packed my bag”. As the agent’s policy refines, the task dynamically expands, serving as an automatic curriculum. This innovative framework enables a thorough evaluation of memory-based agents, emphasizing their overall effectiveness beyond mere interaction efficiency with the environment.
To foster a more inclusive research landscape, we contributed an open-source PPO baseline powered by Transformer-XL (TrXL). This baseline employs an attention mechanism applied to an episodic memory with a sliding window. When benchmarked against Memory Gym, the TrXL baseline competes head-to-head with another prominent baseline rooted in Gated Recurrent Unit (GRU). Notably, in finite environments, TrXL demonstrates superior effectiveness in Mortar Mayhem and enhanced sample efficiency in Mystery Path. However, in the Searing Spotlights environment, GRU emerges as the more efficient architecture.
Perhaps our most unexpected revelation is the comeback of GRU in the endless environments. GRU surpassed TrXL by large margins, while also being computationally more efficient. Further probing into the Endless Mortar Mayhem environment revealed only marginal improvements for TrXL upon enriching the learning signal.
@@ Acknowledgements
Our sincere appreciation goes to Vincent-Pierre Berges for his enlightening discussions that greatly enriched our work. We also extend our gratitude to Andrew Lampinen for his invaluable insights on utilizing Transformer-XL as episodic memory. Special thanks are due to Günter Rudolph for his unwavering support. This project would not have been possible without the generous computing time provided by the Paderborn Center for Parallel Computing (PC2) and the support from the Linux-HPC-Cluster (LiDO3) at TU Dortmund. We are deeply thankful for all of their contributions.
@@ Appendix B Hyperparameters
Number of TrXL Blocks 3 2 4
Block Dimension 384 256 512
@@
Positional Encoding Relative None Learned
@@
Our hyperparameter search was conducted using a customized implementation built upon the optuna tuning framework (Akiba et al., 2019). The objective was to discover a set of hyperparameters that demonstrate strong performance across both our GRU and TrXL baselines and Memory Gym’s environments. Considering limited resources and the computational expense of tuning, we restricted the search space to a limited number of options. The majority of tuning experiments focused on the finite environments. Therefore, we do not claim to provide optimal hyperparameters.
@@
Appendix C GTrXL and LSTM Results on the Finite Environments
@@
Appendix D Wall-Time Efficiency
@@
Figure 11: Wall-time
We further take a look at the wall-time of various agents trained on Mystery Path (Figure D). It becomes apparent that TrXL and GRU are quite on par concerning wall-time efficiency, whereas TrXL needs fewer samples as shown in section 5.1. On first sight, it seems surprising that GTrXL is faster than TrXL (Table 6) even though GTrXL is more complex. This is due to the more expensive resets of the environment. Better policies have shorter episodes and thus reset more frequently impairing wall-time. So as GTrXL’s and LSTM’s policies are inferior, their wall-time is faster. These results were obtained from the LiDo3 high-performance cluster [^2] https://lido.itmc.tu-dortmund.de/. Each run utilized 32 cores (AMD Epyc 7542), an A100 GPU, and 200GB RAM.

### v5

@@ 6 Conclusion
In this study, we advanced Memory Gym’s environments to novel endless tasks, tailored to benchmark memory-based Deep Reinforcement Learning algorithms. These tasks feature a mounting challenge similar to the cumulative memory game “I packed my bag”. As the agent’s policy refines, the task dynamically expands, serving as an automatic curriculum. This innovative framework enables a thorough evaluation of memory-based agents, emphasizing their overall effectiveness beyond mere interaction efficiency with the environment. Our experimental results reveal that agents trained on these endless tasks consistently show stronger capabilities than those trained on finite tasks, demonstrating that finite environments offer only a restricted perspective on an agent’s full memory capabilities.
To foster a more inclusive research landscape, we contributed an open-source PPO baseline powered by Transformer-XL (TrXL). This baseline employs an attention mechanism applied to an episodic memory with a sliding window. When benchmarked against Memory Gym, the TrXL baseline competes head-to-head with another prominent baseline rooted in Gated Recurrent Unit (GRU). Notably, in finite environments, TrXL demonstrates superior effectiveness in Mortar Mayhem and enhanced sample efficiency in Mystery Path. However, in the Searing Spotlights environment, GRU emerges as the more efficient architecture. Perhaps our most unexpected revelation is the comeback of GRU in the endless environments. GRU surpassed TrXL by large margins, while also being computationally more efficient. Further probing into the Endless Mortar Mayhem environment revealed only marginal improvements for TrXL upon enriching the learning signal.
@@ Acknowledgements
Our sincere appreciation goes to Vincent-Pierre Berges and Fabian Ostermann for their enlightening discussions that greatly enriched our work. We also extend our gratitude to Andrew Lampinen for his invaluable insights on utilizing Transformer-XL as episodic memory. Special thanks are due to Günter Rudolph for his unwavering support. This project would not have been possible without the generous computing time provided by the Paderborn Center for Parallel Computing (PC2) and the support from the Linux-HPC-Cluster (LiDO3) at TU Dortmund. We are deeply thankful for all of their contributions.
@@ Appendix B Hyperparameters
Number of TrXL Layers 3 2 4
TrXL Block Dimension 384 256 512
@@
Positional Encoding Absolute None Learned
@@
Our hyperparameter search was conducted using a customized implementation built upon the optuna tuning framework (Akiba et al., 2019). The objective was to discover a set of hyperparameters that demonstrate strong performance across Memory Gym’s environments, for the GRU and TrXL baselines individually. Considering limited resources and the computational expense of tuning, we restricted the search space to a limited number of options. The majority of tuning experiments focused on the finite environments. Therefore, we do not claim to provide optimal hyperparameters.
@@
Appendix C Results of Minor Baselines on the Finite Environments
(a) Mortar Mayhem Act Grid
(b) Mortar Mayhem Grid
(c) Mystery Path Grid (Cues On)
(d) Mystery Path Grid
(e) Searing Spotlights
Figure 10: Performance comparison of several minor baselines on instances of Memory Gym’s finite environments. Mortar Mayhem Act Grid (a) provides commands as a fully observable vector, bypassing the Clue Task. Both (a) and Mortar Mayhem Grid (b) utilize grid-like locomotion. Mystery Path Grid operates on grid-like locomotion, with (d) rendering the the goal and origin to the agent’s observation. (e) hides these. Searing Spotlights (f) is not varied, while the minor baselines leverage observation reconstruction (Obs. Rec.) in this environment.
This set of experiments reaffirms that agents trained on Memory Gym require memory. To show this, we train several naive baselines on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with absolute positional encoding (Vaswani et al., 2017) provided as vector observation. Note that the hyperparameters for these baselines are not tuned.
In Mortar Mayhem Act Grid, where the commands are completely provided as vector observation, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend, while the one with absolute positional encoding follows closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the complete task is present (Figure 10(b)), only GRU and TrXL prove to be effective.
We obtain a similar impression when training on Mystery Path Grid where the origin and goal are not perceivable by the agent (Figure 10(d)). If the origin and the goal are visible, a horizon of 16 frames is sufficient to train an effective policy as shown by the frame stacking agent (Figure 10(c)). Stacking 4 frames or leveraging positional encoding leads to an IQM success rate of 40% with a slight trend upward. The memory-less agent’s success rate goes up to about 11%.
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 10(e). Only in this experiment, the naive baselines also leverage observation reconstruction.
Appendix D GTrXL and LSTM Results on the Finite Environments
@@
Appendix E Wall-Time Efficiency
@@
Figure 12: Wall-time efficiency curves retrieved from Mystery Path
We further take a look at the wall-time of various agents trained on Mystery Path (Figure 12). It becomes apparent that TrXL and GRU are quite on par concerning wall-time efficiency, whereas TrXL needs fewer samples as shown in section 5.3. On first sight, it seems surprising that GTrXL is faster than TrXL (Table 6) even though GTrXL is more complex. This is due to the more expensive resets of the environment. Better policies have shorter episodes and thus reset more frequently impairing wall-time. So as GTrXL’s and LSTM’s policies are inferior, their wall-time is faster. These results were obtained from the LiDo3 high-performance cluster [^2] https://lido.itmc.tu-dortmund.de/. Each run utilized 32 cores (AMD Epyc 7542), an A100 GPU, and 200GB RAM.
@@
Appendix F Simulation Speed of the Environments
Memory Gym’s environments are implemented using Python and PyGame [^3] https://www.pygame.org. The simulation speeds are shown in Table 7. At first glance, the endless variants appear slower than the finite ones. This discrepancy is attributed to the shorter episodes in endless variants, which rely on less effective constant action policies, thereby increasing the impact of environment reset costs. Compared to related memory benchmarks, our environments perform exceptionally well, with only Procgen (Cobbe et al., 2020) being faster (Pleines et al., 2023). The training costs in our experiments are primarily due to the inference time of the baseline models.
Table 7: Comparison of the simulation speed of Memory Gym’s environments. A constant action is executed to measure the speed, and the episode reset duration is included in the steps per second. The measurements were done on an AMD Ryzen 7 2700X, aggregating 1000 episodes per environment.
Environment Mean (steps/sec) Std (steps/sec)
Mystery Path 14033 158
Mortar Mayhem 13187 208
Endless Mortar Mayhem 9330 452
Endless Mystery Path 7278 555
Endless Searing Spotlights 5828 106
Searing Spotlights 5789 165